OpenStudy (anonymous):

Solve. (x-16)-(2x-7)=1

OpenStudy (anonymous):

Distribute the negative to both 2x and -7 which then gives you -2x and positive 7. Then combine like terms (-16 and 7, -2x and x). Get x alone afterwards

OpenStudy (rosedewittbukater):

-x -9 = 1 -x = 10 x = 10

OpenStudy (rosedewittbukater):

I subtracted the like terms (x-2x) = -x and (-16-(-7)) = -16+7 = 9 to get -x -9 =1 then I added 9 to both sides to get x by itself -x = 10 and finally divided both sides by -1 so x = 10

OpenStudy (anonymous):

Thank you